# S2-Diffusion: Generalizing from Instance-level to Category-level Skills in Robot Manipulation

### v2

@@ 13                 $\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(a^{k},k,f_{v},q))$
ServeMug CloseDoor TurnOnMicrowave TurnOffFaucet MoveSoda
BC-RNN 0.00 0.02 0.00 0.26 0.02
BC-Transformer 0.14 0.58 0.78 0.26 0.08
Diffusion Policy 0.12 0.42 0.34 0.32 0.16
S2-Diffusion (Ours) 0.72 0.78 0.84 0.70 0.82
@@ V-C Learning Semantic Diffusion Policy
To effectively utilize our open-vocabulary spatial-semantic representations, we adopt the CNN-based diffusion policy architecture [9] as our decision-making backbone. We use Denoising Diffusion Implicit Models (DDIM) [53] as the noise scheduler. In our paper, we denote the observation with $s_{t}$ that is composed of spatial-semantic feature $f_{v}$ and robot proprioceptive state $q$. The training loss in Equation 2 is defined as:
$\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(\alpha_{k}a^{0}+\beta_{k}\epsilon^{k},k,f_{v},q)),\$ (4)
@@ VI-A Simulation Experiments
Experiment Setup. For our simulation experiments, we take advantage of a recent open-sourced large-scale simulation environment, RoboCasa [52], which provides expert demonstrations for diverse everyday tasks. We evaluate our method and the baselines in a set of single-stage tasks as shown in Fig. 4: ServeMug, CloseDoor, TurnOnMicrowave, TurnOffFaucet, MoveSoda using the provided $50$ expert demonstrations from RoboCasa.
@@
Baselines. We compare our method with three baseline methods in simulation: 1) BC-RNN: a behavior cloning method with recurrent neural network implementation; 2) BC-Transformer: a behavior cloning imitation learning method with transformer architecture as the policy implementation [54]; 3) Diffusion Policy: the image-based diffusion policy of [9]. For real-world experiments, we compare our method against the diffusion policy. In all our experiments, we train for $500$ epochs on an NVIDIA RTX 4090 GPU. We use a batch size of $64$ and a learning rate of $1\text{e}{-4}$.
@@
Simulation Results. We take advantage of the open-sourced implementation for the baseline methods from RoboMimic [54]. In the demonstration datasets for each task, we assume variations set by RoboCasa in the color and shape of the target object, as well as differences in the background environment. To test the performance of our method and the baselines, we evaluate $50$ trials for each task and the corresponding success rates are listed in Table I, demonstrating the superior performance of the proposed S2-Diffusion approach.
@@
Due to the challenging variations in target object and background seen within the RoboCasa dataset, classical imitation learning policies struggle to solve the specific single-stage manipulation task. While the baseline methods—BC-RNN, BC-Transformer, and Diffusion Policy—show relatively poor performance, particularly on tasks like ServeMug, CloseDoor, and MoveSoda, S2-Diffusion consistently achieves the highest success rates. For instance, S2-Diffusion outperforms the other methods with a success rate of 0.72 on ServeMug and 0.78 on CloseDoor, while the best-performing baseline, BC-Transformer, achieves only 0.14 and 0.58, respectively. On other tasks like TurnOnMicrowave and MoveSoda, S2-Diffusion maintains a strong performance with success rates of 0.84 and 0.82. These results highlight the effectiveness of S2-Diffusion in improving task success across a diverse range of manipulation tasks. Similar to findings in RoboCasa [52], the image-based Diffusion Policy is sensitive to in object color and background. In contrast, by leveraging spatial-semantic features our method shows better robustness and generality compared to state-of-the-art alternatives. We attribute this to the rich visual representation of combining semantic information and depth estimation for the workspace scene.
@@ VI-B Real-World Experiments
Experiment setup. We evaluate our S2-Diffusion method on two real-world category tasks: whiteboard-wiping and bowl-to-bowl scooping. We collected $40$ and $60$ expert demonstrations for the red-whiteboard-wiping and rice-bowl-to-bowl scooping tasks instances respectively. The demonstrations were obtained by teleoperating a $7$-DOF Franka Panda manipulator using the Quest2ROS [55] Oculus app. A single camera providing RGB observation was mounted on the end-effector as well as task-specific tools - such as a sponge for the whiteboard wiping and a spoon for the scooping task. The teleoperation setup is shown in Fig. 5. The language prompts for obtaining the semantic observations were "handwriting. sponge." for the whiteboard wiping task and "rice. bowl." for the bowl-to-bowl scooping task.
Figure 5: Real-world setup for the wiping task showing the teleoperation controller held by the operator on the right.
Baselines and Ablations. We compare our S2-Diffusion method with the visuomotor diffusion method from [9] which uses RGB and proprioceptive information as a baseline, as well as perform a number of ablations on the scooping task - namely a version of our method that only has access to semantic observations (Semantic-Diffusion) and one that only observes the spatial observations (Spatial-Diffusion).
@@
Whiteboard-wiping: The task involves wiping scribbles of different colors (red, black, green) from a whiteboard using the end-effector-mounted wet sponge. The results are shown in Fig. 6. We define success as the robot completely removing the targeted handwriting using a sponge within $15$ seconds. For the task instances that are covered by the training data (red-whiteboard-wiping), the baseline is able to succeed in all nine trials, leading to a $100\%$ success rate, however, once the policy is deployed on different instances of the task such as black-whiteboard-wiping the performance deteriorates to $4/9$ ($44\%$) and for the green-whiteboard-wiping task even to $0\%$.
@@
This clearly shows how even small changes in the RGB observations such as changing the color of the scribbles can lead to great deterioration and even complete failure of the skill on this instance - underscoring that the skill learning was indeed on the instance-level and not category-level. On the other hand, our S2-Diffusion method has consistent performances across all tasks. These results highlight the ability of our S2-Diffusion to learn category-level skills for handling novel tasks without requiring additional training or fine-tuning, a capability that is essential for real-world robotic applications.
@@
Bowl-to-bowl-scooping: The scooping task is to scoop granular materials from one bowl to another. The amount of material successfully scooped into the target bowl is measured in grams. Task success is defined as the policy’s ability to scoop at least $3$ grams of material into the target bowl in under $30$ seconds. In Fig. 6 the success rates are reported over $10$ trials. All real-world experimental videos are available on the project website.
For the in-distribution rice-bowl-to-bowl-scooping task, our S2-Diffusion and baseline achieve success rates of $1.0$ and $0.8$, respectively. However, when transferring the policy to three unseen tasks—–choco-cereal-btb-scooping, hearts-cereal-btb-scooping, and mixed-cereal-btb-scooping—–the baseline diffusion policy fails entirely, with $0.0$ success rate, while our S2-Diffusion policy maintains a high success rate of approximately $0.8$ by changing the semantic prompt to "cereal. bowl." for all three instances. This demonstrates that our method effectively generalizes from individual instances to unseen other instances of the same category.
@@
We can see frames of the baseline diffusion policy and our method in Fig. 7, where the baseline can only succeed on the instance-level task it was trained on and fails to scoop out cereal as the visual observations are too different. Our method on the other hand extracts the semantic mask from the prompt as well as the spatial information via the synthetic depth observation. This leads to the successful execution of the scooping task and to the learning of a successful category-level skill trained on individual instances only using a single RGB observation (same as the baseline) as the original input.
@@
Figure 7: Frames of the baseline diffusion policy and our S2-Diffusion method for real-world scooping tasks. We show the corresponding semantic mask and synthetic depth for each frame. The baseline can only succeed on the instance-level task it was trained on. Our method succeeds in learning a category-level skill trained on individual instances only using a single RGB observation.
@@ VI-C Ablation Evaluation
To investigate the significance of integrating spatial-semantic representation, we conduct additional experiments by training the policy using only the semantic representation (Semantic-Diffusion) and only the depth map estimation (Spatial-Diffusion) on the rice and mixed-cereal-bowl-to-bowl-scooping tasks. The results of these experiments are presented in Fig. 8. The full model S2-Diffusion, which integrates both modalities, achieves the highest success rates, significantly outperforming the ablated versions both on the seen task and also the novel task.
When the depth information is excluded, the performance of the policy drops drastically, with success rates falling to $0.4$ and $0.5$ for the seen and novel tasks, respectively. This significant decline underscores the critical role that spatial information plays in successfully executing 3D tasks. In contrast, when the mask (semantic representation) is removed, the performance also declines but to a lesser extent compared to the removal of depth information. While the policy’s ability to generalize across tasks is still reasonable in this case, it clearly benefits from the combined use of both modalities.
From these results, we conclude that both semantic representation and depth estimation are indispensable for learning a generalizable skill trained on individual instances. The integration of spatial-semantic features enhances the robustness of the policy, enabling better performance across both familiar and novel environments, and ultimately leading to more effective task execution in 3D spaces.
Figure 8: Ablation study of Semantic and Spatial-diffusion only on a seen task instance (left) and an unseen task instance (right). We show how the semantic component and spatial component contribute to the performance of the policy. S2-Diffusion outperforms models trained with only semantic or spatial observations.
@@ VII Limitations
As Fig. 2 shows, the functional goal of flipping or scooping is discretized in different category-level skills such as bowl-to-bowl and pile-to-container scooping. While our method does generalize from instance-level tasks to category-level tasks, in its current form it is not able to generalize well from one instance-level skill to an instance of a different category task. i.e. when only training on rice-bowl-to-bowl the performance will degrade on a sand-pile-to-container task.
While Fig. 2 makes the ordering of skills into instance and category levels look straightforward, there are some corner cases that would need to be addressed if this view was expanded into a more complex ontology. For one, the functional goal can depend on more context than what is shown in our example. For instance, flipping is presented in our example as a way to turn over an object, but it could also be interpreted as an action the robot should perform with its own body. Furthermore, given how current learning-from-demonstration techniques are implemented, the category bowl-to-bowl scooping might require additional qualifiers to accurately encompass the correct instances. For example, if we were to generalize this skill to ice cream bowl-to-bowl scooping, the policy would fail, as the forces required are completely different from those observed in the training instances. Ice cream, especially when frozen, requires significantly more force to be scooped in the first place. However, as the ice cream begins to melt, and the required forces become more similar to those seen during training, there is a point at which our category skill bowl-to-bowl scooping would likely work. We leave exploring such nuances to future work.
Another limitation of our work is its dependency on the performance of pre-trained vision foundation models. Naturally if the semantic or spatial estimate is poor, the resulting category-level skill may generalize poorly. To address this, we plan to investigate fine-tuning VLMs in conjunction with the downstream robot policy network in future research.
@@ VIII CONCLUSION
In this work, we have introduced S2-Diffusion, an open-vocabulary spatial-semantic diffusion policy that enables generalization from instance-level training data to category-level skills in robotic manipulation. Our method integrates semantic understanding with spatial representations, leveraging vision foundation models to learn a policy that is invariant to task-irrelevant visual changes and generalizes across different instances of the same category. This allows robots to transfer learned skills beyond their training data without requiring additional fine-tuning or retraining.
Through extensive simulations and real-world evaluations, we demonstrated that S2-Diffusion outperforms the baselines. In particular we showed that:
  • Spatial-Semantic representations enhance generalization – By combining spatial information with semantic segmentation, our method enables robots to focus on task-relevant features, improving instance-to-category transfer.
  • Efficient real-time execution using only a single RGB camera – Unlike approaches that require multi-view setups or additional depth sensors, S2-Diffusion extracts meaningful scene representations from a single RGB image, making it practical for real-world deployment.
  • Category-level generalization – Our evaluations on a number of robotic manipulation tasks, including real-world wiping and scooping, demonstrate that the method successfully generalizes across unseen instances within the same category, achieving high performance where baseline policies fail.
In summary, S2-Diffusion policies are less constrained by their training distribution and constitute a step toward the ability of humans to generalize skills across variations in objects, materials, and environments instances.

### v3

@@ V-B Open-Vocabulary Spatial-Semantic Representation
ServeMug CloseDoor TurnOnMicrowave TurnOffFaucet MoveSoda TurnOnStove
BC-RNN 0.02 [0.01, 0.07] 0.03 [0.01, 0.08] 0.05 [0.02, 0.11] 0.27 [0.19, 0.36] 0.06 [0.03, 0.12] 0.26 [0.18, 0.35]
BC-Transformer 0.16 [0.10, 0.24] 0.55 [0.45, 0.64] 0.70 [0.60, 0.78] 0.31 [0.23, 0.41] 0.10 [0.06, 0.17] 0.34 [0.25, 0.44]
Diffusion Policy 0.15 [0.09, 0.23] 0.52 [0.42, 0.62] 0.41 [0.32, 0.51] 0.38 [0.29, 0.48] 0.22 [0.15, 0.31] 0.50 [0.40, 0.60]
S2-BC-Transformer 0.54 [0.44, 0.63] 0.77 [0.68, 0.84] 0.72 [0.63, 0.80] 0.55 [0.45, 0.64] 0.44 [0.35, 0.54] 0.60 [0.50, 0.69]
S2-Diffusion (Ours) 0.75 [0.66, 0.82] 0.80 [0.71, 0.87] 0.80 [0.71, 0.87] 0.77 [0.68, 0.84] 0.83 [0.74, 0.89] 0.88 [0.80, 0.93]
TABLE II: Simulation Experiments in SAPIEN
Method InsertPencil HangMug
Seen Unseen Seen Unseen
GenDP 0.95 [0.76, 0.99] 0.85 [0.64, 0.95] 0.90 [0.70, 0.97] 0.80 [0.58, 0.92]
S2-Diffusion (Ours) 0.90 [0.70, 0.97] 0.90 [0.70, 0.97] 0.85 [0.64, 0.95] 0.90 [0.70, 0.97]
@@ V-C Learning Semantic Diffusion Policy
To effectively utilize our open-vocabulary spatial-semantic representations, we adopt the CNN-based diffusion policy architecture [2] as our decision-making backbone. We use Denoising Diffusion Implicit Models (DDIM) [42] as the noise scheduler. In our paper, the observation is composed of spatial-semantic feature $f_{v}$ and robot proprioceptive state $q$. The training loss in Equation 2 is defined as:
$\mathcal{L}=\textit{MSE}(a^{0},\pi_{\theta}(\alpha_{k}a^{0}+\beta_{k}\epsilon^{k},k,f_{v},q)),\\$ (4)
@@ VI-A Simulation Experiments
Experiment Setup. For our simulation experiments, we take advantage of a recent open-sourced large-scale simulation environment, RoboCasa [41], which provides expert demonstrations for diverse everyday tasks. We evaluate our method and the baselines on one task of each atomic task category except navigation and group doors and drawers due to their similarity as shown in Fig. 4: ServeMug, CloseDoor, TurnOnMicrowave, TurnOffFaucet, MoveSoda, and TurnOnStove using the provided $50$ expert demonstrations from RoboCasa. The prompts used were: "Mug", "Door", "Microwave button", "Faucet", "Soda can", "Knob" respectively.
@@
Baselines. We compare our method with three baseline methods in simulation: 1) BC-RNN: a behavior cloning method with recurrent neural network implementation; 2) BC-Transformer: a common behavior cloning method with transformer architecture [43]; 3) Diffusion Policy: the image-based diffusion policy of [2]; for completeness we also compare against 4) S2-BC-Transformer: BC-Transformer trained with spatial-semantic data; and 5) GenDP: a multi-view RGB-D method [22]. As RoboCasa does not provide depth, we compare against GenDP on InsertPencil and HangMug in SAPIEN  [44]. In all our experiments, we train for $500$ epochs on an NVIDIA RTX 4090 GPU. We set the initial learning rate to $1\text{e}{-4}$, applied a $500$-step linear warm-up, and then used a cosine learning rate scheduler for the remainder of training.
@@
Simulation Results. For simulation experiments in RoboCasa, we take advantage of the open-sourced implementation for the baseline methods from RoboMimic [43]. In the demonstration datasets for each task, we assume variations set by RoboCasa in the color of the target object, as well as differences in the background environment. We evaluate $100$ trials with the default seed for each task and the corresponding success rates and the $95\%$ confidence intervals (computed via the Wilson score method for binomial proportions) are listed in Table I. The results are demonstrating matching or statistically significantly superior performance of our S2-Diffusion approach.
@@
Due to the challenging variations in target object and background seen within the RoboCasa dataset, classical imitation learning policies struggle to solve the specific single-stage manipulation task. While the baselines—BC-RNN, BC-Transformer, and Diffusion Policy—show relatively poor performance, particularly on tasks like ServeMug and MoveSoda, S2-Diffusion consistently achieves the highest success rates. This lower performance of both BC and diffusion policy on RoboCasa is likely due to the high diversity of scenes in the expert training data. We also trained S2-BC-Transformer on the spatial-semantic dataset. We extended BC-Transformer because in our prior evaluation it exhibited comparable performance with DP. While S2-BC-Transformer achieves competitive results on some tasks, S2-diffusion demonstrates more robust performance. Additionally, S2-BC-Transformer outperforms the RGB-only version, highlighting the value of spatial-semantic representations for improving the generalization of imitation policies. Similar to findings in RoboCasa [41], the image-based DP is sensitive to in object color and background. In contrast, by leveraging spatial-semantic features our method shows better robustness and generality compared to state-of-the-art alternatives. We attribute this to the rich visual representation of combining semantic information and depth estimation for the workspace scene. Finally, Table II shows that our RGB method performs on par with the multi-view RGB-D method GenDP.
@@ VI-B Real-World Experiments
Experiment setup. We evaluate our S2-Diffusion method on two real-world category tasks: whiteboard-wiping and bowl-to-bowl scooping. We collected $40$ and $60$ expert demonstrations for the red-whiteboard-wiping and rice-bowl-to-bowl scooping tasks instances respectively. The demonstrations were obtained by teleoperating a $7$-DOF Franka Panda manipulator using the Quest2ROS [45] Oculus app. A single camera providing RGB observation was mounted on the end-effector as well as task-specific tools - such as a sponge for the whiteboard wiping and a spoon for the scooping task. The language prompts for obtaining the semantic observations were "handwriting. sponge." for the whiteboard wiping task and "rice. bowl." for the bowl-to-bowl scooping task.
Baselines and Ablations. We compare our S2-Diffusion method with the visuomotor diffusion method from [2] which uses RGB and proprioceptive information as a baseline, as well as perform a number of ablations on the scooping task - namely a version of our method that only has access to semantic observations (Semantic-Diffusion) and one that only observes the spatial observations (Spatial-Diffusion). Finally, we evaluate generalization to transparent objects.
@@
Whiteboard-wiping. The task involves wiping scribbles of different colors (red, black, green) from a whiteboard using the end-effector-mounted wet sponge. The results are shown in Fig. 5 including the $95\%$ Wilson confidence interval. We define success as the robot completely removing the targeted handwriting using a sponge within $15$ seconds. For the task instances that are covered by the training data (red-whiteboard-wiping), the baseline is able to succeed in all nine trials, leading to a $100\%$ success rate, however, once the policy is deployed on different instances of the task such as black-whiteboard-wiping the performance deteriorates to $4/9$ ($44\%$) and for the green-whiteboard-wiping task even to $0\%$. This clearly shows how even small changes in the RGB observations such as changing the color of the scribbles can lead to great deterioration and even complete failure of the skill on this instance - underscoring that the skill learning was indeed on the instance-level and not category-level. On the other hand, our S2-Diffusion method has consistent performances across all tasks. These results highlight the ability of our S2-Diffusion to learn category-level skills for handling novel tasks without requiring additional training or fine-tuning, a capability that is essential for real-world robotic applications. Note that the results of black-whiteboard-wiping are not statistically significant - we attribute this to the fact that the black observations are more similar to red than the green scribbles.
@@
Bowl-to-bowl-scooping. The scooping task is to scoop granular materials from one bowl to another. The amount of material successfully scooped into the target bowl is measured in grams. Task success is defined as the policy’s ability to scoop at least $3$ grams of material into the target bowl in under $30$ seconds. In Fig. 5 the success rates are reported over $10$ trials. All real-world experimental videos are available on the project website.
@@
For the in-distribution rice-bowl-to-bowl-scooping task, our S2-Diffusion and baseline achieve success rates of $1.0$ and $0.8$, respectively. However, when transferring the policy to three unseen tasks—–choco-cereal-btb-scooping, hearts-cereal-btb-scooping, and mixed-cereal-btb-scooping—–the baseline diffusion policy fails entirely, with $0.0$ success rate, while our S2-Diffusion policy maintains a high success rate of approximately $0.8$ by changing the semantic prompt to "cereal. bowl." for all three instances. This demonstrates that our method effectively generalizes from individual instances to unseen other instances of the same category. We can see frames of the baseline diffusion policy and our method in Fig. 6, where the baseline can only succeed on the instance-level task it was trained on and fails to scoop out cereal as the visual observations are too different. Our method on the other hand extracts the semantic mask from the prompt as well as the spatial information via the synthetic depth observation. This leads to the successful execution of the scooping task and to the learning of a successful category-level skill trained on individual instances only using a single RGB observation (same as the baseline) as the original input.
@@
Figure 6: Frames of the baseline diffusion policy and our S2-Diffusion method for real-world scooping tasks. We show the corresponding semantic mask and synthetic depth for each frame. The baseline can only succeed on the instance-level task it was trained on. Our method succeeds in learning a category-level skill trained on individual instances.
@@ VI-C Ablation Evaluation
To investigate the significance of integrating spatial-semantic representation, we conduct additional experiments by training the policy using only the semantic representation (Semantic-Diffusion) and only the depth map estimation (Spatial-Diffusion) on the rice and mixed-cereal-bowl-to-bowl-scooping tasks. The results of these experiments are presented in Fig. 7. The full model S2-Diffusion, which integrates both modalities, achieves the highest success rates, significantly outperforming the Semantic-Diffusion and also etching out the Spatial-Diffusion versions both on the seen task and also the novel task. When the depth information is excluded, the performance of the policy drops drastically, with success rates falling to $0.4$ and $0.5$ for the seen and novel tasks, respectively. This significant decline underscores the critical role that spatial information plays in successfully executing 3D tasks. To evaluate generalization to transparent objects, we add a task where it has to close a transparent container. For reference, we also compare against a DP with depth camera input, trained directly on the transparent container. We find that S2 Diffusion only has a small reduction in success rate while the depth camera policy was not able to complete the task - illustrating the robustness of the S2 approach in common kitchen tasks involving transparent or reflective surfaces - where depth sensors struggle. In contrast, when the mask (semantic representation) is removed, the performance also declines, but to a lesser extent compared to the removal of depth information. While the policy’s ability to generalize across tasks is still reasonable in this case, it benefits from the combined use of both modalities.
Figure 7: Ablation study of Semantic and Spatial-diffusion only on a seen task instance (left) and an unseen task instance (right). S2-Diffusion outperforms models trained with only semantic or spatial observations.
@@ VII Limitations
As Fig. 2 illustrates, functional goals like flipping or scooping can be discretized into different category-level skills such as bowl-to-bowl and pile-to-container scooping. While our method generalizes from instance-level to category-level tasks, it does not generalize well across different categories; for example, a policy trained on rice-bowl-to-bowl scooping struggles with a sand-pile-to-container task. Although the abstraction hierarchy in Fig. 2 appears straightforward, real-world functional goals can depend on more complex context and categories like bowl-to-bowl scooping may require additional qualifiers. For example, transferring frozen ice cream differs significantly from transferring cereals due to the forces required, although generalization may improve as the ice cream melts. Addressing such nuanced skill ontologies is an avenue for future work. Another limitation is the dependency on the performance of pretrained VLMs as well as proper text prompts for accurate semantic segmentation. Poor semantic or spatial estimates can degrade category-level skill generalization, particularly for greatly different object instances. Fine-tuning VLMs jointly with the policy network is a promising direction for mitigating this dependency.
@@ VIII CONCLUSION
We introduced S2-Diffusion, an open-vocabulary spatial-semantic diffusion policy that enables generalization from instance-level training data to category-level skills in robotic manipulation. By integrating semantic understanding and spatial representations via vision foundation models, our method learns policies invariant to task-irrelevant visual changes, allowing skill transfer without additional fine-tuning. Through extensive simulations and real-world evaluations, we demonstrated that S2-Diffusion outperforms baselines. In particular:
  • Spatial-Semantic representations enhance generalization: By combining spatial and semantic information, our method focuses on task-relevant features, enabling instance-to-category transfer.
  • Efficient real-time execution: S2-Diffusion requires only a single RGB camera, avoiding multi-view setups or depth sensors, making it practical for real-world deployment and attractive for mobile settings.
  • Category-level generalization: Our evaluations show strong generalization across unseen instances, achieving high performance where baseline policies fail.
Overall, S2-Diffusion represents a step toward enabling robots to generalize skills across variations in objects, materials, and environments, similar to human capabilities.